OpenStudy (anonymous):

Points Y and Z are the same distance from point X on a number line. The coordinate of point X is –7 and the coordinate of point Y is –1. What is the coordinate of point Z?

OpenStudy (anonymous):

hello again

OpenStudy (stefrheart):

Hi :) So im guessing that Y and Z are a line and that X is the midpoint

OpenStudy (anonymous):

i guess so

OpenStudy (stefrheart):

Lol jk this is actually easier than it seems First find the distance between X and Y

OpenStudy (anonymous):

ok

OpenStudy (stefrheart):

So what is it?

OpenStudy (anonymous):

ok they are the same y=-1 so z has to equal 1

OpenStudy (anonymous):

i think im wrong

OpenStudy (stefrheart):

No find the distance from -7 to -1 So it you were to count from a number line you would count how many spaces

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

6 units away

OpenStudy (anonymous):

is that right

OpenStudy (stefrheart):

yes now add 6 to both X and Y

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

it -12

OpenStudy (stefrheart):

-12?

OpenStudy (anonymous):

no -2

OpenStudy (stefrheart):

Shouldnt it be -13 and/or 5

OpenStudy (anonymous):

no you said to add 6+ y=-1+ X=-7 and that equal -2

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

is anyone there

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

i cant take this anymore bye

OpenStudy (texaschic101):

-7 - (-1) = -7 + 1 = 6.....6 spaces between X and Y. So if Y is 6 spaces from X, then Z is also 6 spaces from X.  I can't draw on this thing...lol X = -7 Y = -1 Z = -13